An 8.5x10^2 kg elevator is pulled up 32.0m at a constant velocity of 1.4m/s. Calculate the power output of the motor.
1
Expert's answer
2019-07-29T08:24:57-0400
"{\rm{Power}}=\frac{\rm{Work\;done}}{\rm{Time}}=\frac{mgh}{t}=mgv"

"=8.5\times 10^2\:\rm{kg}\times 9.8\:\rm{N/kg}\times 1.4\:\rm{m/s}=11662\:\rm{W}"